You have five 10.0 F capacitors. Show all possible connections of all five capacitors to produce an equivalent capacitance of


a. 50.0 F

b. .5 F

c. 35 F​


1
Expert's answer
2022-03-23T07:15:57-0400

a)




C =C1+C2+C3+C4+C5 = 10 + 10 + 10 +10 +10 = 50 F

b)


"C=\frac{C1(\frac{C2C3}{C2+C3}+\frac{C4C5}{C4+C5})}{C1+(\frac{C2C3}{C2+C3}+\frac{C4C5}{C4+C5})}=\frac{10(\frac{10\cdot 10}{10+10}+\frac{10\cdot 10}{10+10})}{10+(\frac{10\cdot 10}{10+10}+\frac{10\cdot 10}{10+10})}=5\space F"


c)

"C=C1+C2+C3+\frac{C4C5}{C4+C5}=10+10+10+\frac{10\cdot 10}{10+10}=35\space F"